OpenStudy (xbeenchantedx):

how do you solve this linear eq. by graphing: eq 1: y= x+ 4 eq 2: 2x - 3y= -7 i know that the ordered pair that they both intersect is (-5, -1).

OpenStudy (anonymous):

Where both lines intersect is the solution to the system of equations. y = -1; x = -5

OpenStudy (xbeenchantedx):

i know but i dont know how to change both equations to y=mx+b form in order to graph it

OpenStudy (anonymous):

We can use our algebra. For the second equation, \[\begin{matrix}2x - 3y = -7 \\ -3y = -2x -7 \\ 3y = 2x + 7 \\ y = {2 \over 3} x + {7 \over 3}\end{matrix}\]

OpenStudy (xbeenchantedx):

it comes out to this fraction but i havent really learned how to find the yint as a fraction and to graph it as a fraction